The basement of a building is flooded with water to a depth of 18 inches. The basement is 52 feet long and 38 feet wide. Two pumps each pumping at the rate of 34 gallons per minute are used to drain the water. How many hours does it take to completely drain the water?
I’m not sure how to even start the problem. Do I need to find the volume? What formula do I use?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

Volume of water is 52*38*1.5 until feet (1.5 feet is 18 inches). V=length*width*height
That product is 2964 ft^3.
1 cubic foot of water=7.48 gallons
so the volume is 2964 ft^3*7.48 gallons/cubic foot=22170.72 gallons,
Now divide that by 2040 gallons per hour (34 gallons/minute * 60 minutes)
=10.87 hours.

Answer by ikleyn(52781)

.

            In his calculations,  tutor  @Boreal accounted for one pump instead of two.
            I came to fix this deficiency.

Volume of water is 52*38*1.5 cubic feet (1.5 feet is 18 inches). V = length*width*height


That product is 2964 ft^3.


1 cubic foot of water=7.48 gallons
so the volume is 2964 ft^3*7.48 gallons/cubic foot=22170.72 gallons.


Now divide that by 8080 gallons per hour (34*2 = 68 gallons/minute * 60 minutes)
=5.434 hours.

Solved  (in a right way).